A freight train passes through a 750-m long tunnel at 50km/h. The back of the train exits the tunnel 1.5 min after the front of the train enters it. What is the length of the train, in meters?
1
Expert's answer
2020-03-19T13:09:43-0400

Let length of train=x metres

Time taken=1.5 min to cover distance=(x+750) metres

Speed of train=50km/hr

=(5000/6) m/min

Therefore (5000/6)=(x+750)/1.5

x=1250-750

=500 metres

Length of train is 500 metres